OpenStudy (anonymous):

The slope of a line is 0, and the y-intercept is 6. What is the equation of the line written in slope-intercept form? A.) y = 6 B.) y = 6x C.) y = x + 6

OpenStudy (johnweldon1993):

\[y = mx + b\] 'm' represents the slope 'b' represents the y-intercept so plug in your slope (0) and your y-intercept (6) and what does your equation look like?

OpenStudy (anonymous):

C

OpenStudy (johnweldon1993):

not quite...remember anything times (0) = 0 so y = mx + b y = 0(x) + 6 y = 0 + 6 y = 6 make sense?

OpenStudy (anonymous):

yea ok i see

OpenStudy (anonymous):

The slope of a line is 1, and the y-intercept is -1. What is the equation of the line written in slope-intercept form?

OpenStudy (anonymous):

@johnweldon1993